Prove ∇^2{FG} = F ∇^2 F + 2∇F•∇G




1
Expert's answer
2021-09-01T18:02:24-0400

"∇^2(FG)\\
=∇.∇(FG)\\
=∇.[F∇G+G∇F]\\
=∇F.∇G+F∇^2G+∇G.∇F+G∇^2F\\
=F∇G+G∇F+2∇F.∇G\\
\text{NOTE:- I correct the question.}"